Section 05  Differential Geometry and Global Analysis

Section 05: Differential Geometry and Global Analysis

Invited Lectures

Wednesday, 19 August  MA 001

14:00 – 14:45 Leonid Polterovich, Tel Aviv University, Tel Aviv, Israel Geometry on the Group of Hamiltonian Diffeomorphisms (I: p.77) Chair: Helmut H.W. Hofer

15:00 – 15:45 Yongbin Ruan, University of Wisconsin-Madison, Madison, USA Quantum cohomology and its applications (I: p.77) Chair: Helmut H.W. Hofer

Thursday, 20 August  MA 001

16:15 – 17:00 Joachim Lohkamp, Universität Augsburg, Augsburg, Germany Curvature Contents of Geometric Spaces (I: p.75) Chair: Jean-Pierre Bourguignon

17:15 – 18:00 Yakov Eliashberg, Stanford University, Stanford, USA Invariants in Contact Topology (I: p.70) Chair: Jean-Pierre Bourguignon

Friday, 21 August  MA 001

14:00 – 14:45 Simon K. Donaldson, Stanford University, Stanford, USA Positive line bundles in symplectic geometry (I: p.69) Chair: Ulrich Pinkall

15:00 – 15:45 François Labourie, Université Paris-Sud, Orsay, France Large Group Actions on Manifolds (I: p.73) Chair: Ulrich Pinkall

Saturday, 22 August  MA 001

16:15 – 17:00 Dominic Joyce, Lincoln College, Oxford, UK Compact Riemannian manifolds with exceptional holonomy (I: p.73) Chair: Boris Dubrovin

17:15 – 18:00 Sylvestre Gallot, U.M.P.A., École Normale Supérieure de Lyon, Lyon, France Curvature-decreasing maps are volume-decreasing (I: p.71) Chair: Boris Dubrovin

Monday, 24 August  MA 001

14:00 – 14:45 Ulrich Pinkall, TU Berlin, Berlin, Germany Quaternionic line bundles and the differential geometry of surfaces (I: p.75) Chair: Dirk Ferus

15:00 – 15:45 Tobias Holck Colding, Courant Institute, New York University, New York, USA Manifolds and spaces with Ricci curvature bounds (I: p.68) Chair: Dirk Ferus

Differential Geometry and Global Analysis  Section 05

Tuesday, 25 August  MA 001

16:15 – 17:00 Boris Dubrovin, International School for Advanced Studies, Trieste, Italy and Steklov Math. Institute, Moscow, Russia Geometry and analytic theory of Frobenius manifolds (I: p.69) Chair: Alexander I. Bobenko

Wednesday, 26 August  MA 001

14:00 – 14:45 Dmitri Burago, Pennsylvania State University, University Park, USA Hard Balls Gas and Alexandrov Spaces of Curvature Bounded Above (I: p.68) Chair: Frank Duzaar

15:00 – 15:45 Gerhard Huisken, Universität Tübingen, Tübingen, Germany Evolving hypersurfaces by their curvature (I: p.72) Chair: Frank Duzaar

Short Communications and Poster Sessions

Short Communications  Wednesday, 19 August
Kähler Manifolds  MA 041

Chair: Bill Watson

16:15 – 16:30 Keizo Hasegawa, Niigata University, Niigata, Japan A class of compact Kählerian solvmanifolds and a general conjecture (p.71)

16:30 – 16:45 Toyoko Kashiwada, Saitama College, Tokyo, Japan The integrability of almost generalized Hopf manifolds (p.72)

16:45 – 17:00 Bill Watson, St. John’s University, Jamaica, USA The Goldberg Conjecture is true for the four-dimensional total space of an almost Hermitian submersion (p.84)

Topology of Manifolds  MA 041

Chair: Sebastian Goette

17:15 – 17:30 Christina Sormani, Johns Hopkins University, Baltimore, Maryland, USA The Topology of Noncompact Manifolds with Positive Ricci Curvature (p.80)

17:30 – 17:45 Gerald M. Lodder, New Mexico State University, Las Cruces, NM, USA Leibniz Cohomology for Differentiable Manifolds (p.75)

17:45 – 18:00 Sebastian Goette, Universität Freiburg, Freiburg, Germany Equivariant η-invariants and η-forms (p.70)

Miscellaneous Topics  MA 043

Chair: Zheng-Chao Han

17:30 – 17:45 Margaret F. Symington, University of Texas at Austin, Austin, TX, USA Symplectic rational blowdowns (p.81)

17:45 – 18:00 Zheng-Chao Han, Rutgers University, Piscataway, NJ, USA New existence results on conformal metrics with constant interior scalar curvature and constant boundary mean curvature (p.71)

Short Communications  Thursday, 20 August
Special Submanifolds I  MA 041

Chair: Takashi Okayasu

14:00 – 14:15 Vladimir G. Tkachev, Volgograd State University, Volgrograd, Russia Upper estimates of volume growth for minimal quasigraphs (p.82)

14:15 – 14:30 Liu Ximin, Dalian University of Technology, Dalian, P.R. China Positively curved totally complex submanifolds in the Cayley projective space (p.75)

14:30 – 14:45 Takashi Okayasu, Toyama University, Toyama, Japan A Bernstein result for minimal graphs of higher codimension (p.77)

Special Submanifolds II  MA 041

Chair: Ivko M. Dimitric

15:00 – 15:15 Valentin Gabriel Cristea, Technical University of Budapest, Budapest, Hungary Totally umbilical graded submanifolds (p.66)

15:15 – 15:30 S.M.B. Kashani, Tarbiat Madarres University, Tehran, Iran Cohomogeneity One Revolution Hypersurfaces of the Euclidean Space (p.72)

15:30 – 15:45 Ivko M. Dimitric, Pennsylvania State University, Uniontown PA, USA Low-type submanifolds of non-flat complex space forms (p.67)

Short Communications  Saturday, 22 August
Surface Theory  MA 041

Chair: John M. Sullivan

14:00 – 14:15 Udo J. Hertrich-Jeromin, TU, Berlin, Germany Transformations of Isothermic surfaces (p.71)

14:15 – 14:30 Bernd Ammann, Math. Institut der Uni Freiburg, Freiburg, Germany New estimates for the Laplace and Dirac spectrum of 2-Tori and applications to the Willmore functional (p.64)

14:30 – 14:45 John M. Sullivan, University of Illinois, Urbana, IL, USA Triunduloids: Embedded Constant-Mean-Curvature Surfaces with Three Ends (p.81)

Curves  MA 041

Chair: Joachim Giesen

15:15 – 15:30 Kazuyuki Enomoto, Science University of Tokyo, Oshamambe, Hokkaido, Japan Total absolute curvature of nonclosed curves (p.67)

15:30 – 15:45 Joachim Giesen, ETH Zürich, Zürich, Switzerland Geometry of One-Manifolds (p.69)

Short Communications  Monday, 24 August
Integrable Geodesic Flows and Perturbation Problems  MA 041

Chair: Vladimir S. Matveev

16:15 – 16:30 Masayuki Igarashi, Science University of Tokyo, Yamakoshi-gun, Japan On Hermitian-Liouville structures on compact Hermitian surfaces (p.71)

16:30 – 16:45 Li Ma, Tsinghua University, Beijing, China, Peoples Republic of Stable critical point Theory in perturbation problems (p.75)

16:45 – 17:00 Vladimir S. Matveev, Max-Planck-Institut für Mathematik, Bonn, Germany Geodesic equivalence and integrability (p.76)

Local Differential Geometry  MA 043

Chair: Jose Miguel Veloso

16:15 – 16:30 Nikolaus K. Stephanidis, Aristoteles University, Thessaloniki, Greece Über einige Invarianten in der Theorie der Strahlensysteme (p.81)

16:30 – 16:45 Katerina Polyakova, Kaliningrad State University, Kaliningrad, Russia Parallel displacements along manifold of osculating planes of a surface (p.78)

16:45 – 17:00 Jose Miguel Veloso, Universidade Federal Do Para, Belem-PA, Brazil Almost pseudo-Hermitian manifolds (p.83)

Homogeneous Spaces  MA 041

Chair: Friedbert Prüfer

17:15 – 17:30 Takashi Koda, Toyama University, Toyama, Japan Weakly symmetric property of Riemannian 3-symmetric spaces (p.72)

17:30 – 17:45 Jan Slovák, Department of Algebra and Geometry, Masaryk Univesity, Brno, Czech Republic Parabolic Geometries (p.80)

17:45 – 18:00 Friedbert Prüfer, Universität Leipzig, Leipzig, Germany Local homogeneity and differential operators (p.79)

Special Submanifolds/Exterior Differentiation  MA 043

Chair: Peter Rögen

17:15 – 17:30 Iaroslav A. Kopylov, Sobolev Institute of Mathematics, Novosibirsk, Russia Normal solvability of the operator of exterior derivation on some surfaces in ℝ^{N} (p.73)

17:30 – 17:45 Yoshio Matsuyama, Department of Mathematics, Chuo Univ., Tokyo, Japan On totally real submanifolds in CP^{n}(c) (p.76)

17:45 – 18:00 Peter Rögen, Technical University of Denmark, Lyngby, Denmark Embedding and knotting of flat surfaces in 3-space (p.79)

Short Communications  Tuesday, 25 August
Affine Differential Geometry  MA 041

Chair: Salvador D.R. Gigena

14:00 – 14:15 Kurt Leichtweiß, Professor Emeritus, Universität Stuttgart, Stuttgart, Germany On the infinitesimal affine rigidity of ellipsoids in the n-space (p.74)

14:15 – 14:30 Christine Scharlach, FB Mathematik, TU, Berlin, Germany Affine Hyperspheres (p.79)

14:30 – 14:45 Salvador D.R. Gigena, Universidad Nacional de Rosario, Rosario, Argentina On Affine Hypersurfaces with Constant Sectional Curvature (p.70)

General Geometric Structures on Manifolds  MA 043

Chair: Lutz Weimann

14:00 – 14:15 Sergey V. Zuev, Penza State Pedagogical University, Penza, Russia Analytic continuations of SU(2n)-invariant Ricci-flat Kähler metric

14:15 – 14:30 Claudio Gorodski, University of São Paulo, São Paulo, Brazil Contact Sub-Riemannian Symmetric Spaces (p.70)

14:30 – 14:45 Artur E. Fischer, University of California, Santa Cruz CA, USA The Adjoint Formulation and Symplectic Reduction (p.68)

Geodesics and Distance  MA 041

Chair: Jin-ichi Itoh

15:00 – 15:15 Minoru M.T. Tanaka, Tokai University, Hiratsuka, Japan Sard Theorem for the distance function (p.82)

15:15 – 15:30 Lilia Del Riego, Facultad de Ciencias, UASLP, San Luis Potosí, México Exponential map for general sprays and connections (p.66)

15:30 – 15:45 Jin-ichi Itoh, Kumamoto University, Kumamoto, Japan The Lipschitz continuity of the distance function to the cut locus (p.72)

Lorentz Manifolds  MA 043

Chair: Franki Dillen

15:00 – 15:15 Andrzej Krolak, Max-Planck-Institute for Gravitational Physics, Potsdam, Germany Cauchy Horizon Endpoints and Differentiability (p.74)

15:15 – 15:30 Angel L. Garrido Bullon, Departamento de Matematicas, Alcorcon (Madrid), Spain Foliations of Lorentz manifolds (p.66)

15:30 – 15:45 Franki Dillen, Kuleuven, Heverlee, Belgium Ruled Weingarten surfaces in Minkowski space (p.66)

Short Communications  Wednesday, 26 August
Analysis on Manifolds  MA 041

Chair: Marian Gidea

16:15 – 16:30 Yuri G. Borisovich, Voronezh State University, Voronezh, Russia Methods of global analysis in problems of solvability of nonlinear equations and inclusions (p.65)

16:30 – 16:45 Sylvie Paycha, Université Blaise Pascal, Aubiere, France Regularization as a Geometric and Stochastic Tool in Infinite Dimensions (p.78)

16:45 – 17:00 Marian Gidea, Loyola University of Chicago, Chicago, U.S.A. Non-Smooth Dynamical Systems that Exhibit Hyperbolic Behavior (p.69)

Polyhedra/Webs  MA 043

Chair: Hakki Ismail Erdogan

16:30 – 16:45 Anatoly D. Milka, Institute for Low Temperature Physics, Kharkov, Ukraine Linear isometric deformations of polyhedra (p.77)

16:45 – 17:00 Hakki Ismail Erdogan, Istanbul Technical University, Istanbul, Turkey On a class of hexagonal-surface-web of spheres (p.68)

Differential Equations on Manifolds  MA 041

Chair: Congming Li

17:15 – 17:30 Atsushi Tachikawa, Science University of Tokyo, Chiba-ken, Japan A Nonexistence Result of Harmonic Maps between Hadamard Manifolds (p.82)

17:30 – 17:45 Asya Vasilyevna Aminova, Kazan State University, Kazan, Russia The pseudo-Riemannian manifolds with common geodesics (p.64)

17:45 – 18:00 Congming Li, Dept. of Applied Math., Univ. of Colorado, Boulder CO, USA Prescribing Scalar Curvature on S^{n} (p.74)

Poster Session  Friday, 21 August 16:15 – 18:00  MA 144

Philip P. Boalch, Wadham College, Oxford, Great Britain

Symplectic geometry and isomonodromic deformations (p.65) Ilia A. Bogaevski, Independent University of Moscow, Moscow, Russian Federation

Singularities of convex hulls of closed hypersurfaces (p.65) Cristina Draper, Universidad de la Rioja, Lagrono, Spain

Algebras related to Lie triple systems and applications to differential geometry (p.67) Birte Feix, University of Cambridge, Cambridge, Great Britain

Hyperkähler metrics on cotangent bundles (p.68) Lew Friedland, State University of New York, Geneseo, NY, USA

Spectral geometry on almost Hermitian manifolds (p.69) Vladimir A. Geyler, Mordovian State University, Saransk, Russia

Periodic point interactions on a Riemannian manifold (p.251) Valery Konnov, Samara State Pedagogical University, Samara, Russia

Differential geometry of algebraic manifolds (p.73) Alexei G. Kovalev, University of Edinburgh, Edinburgh, United Kingdom

Gluing theorems for anti-self-dual metrics (p.73)

Vasily Olegovich Manturov, Moscow State University, Moscow, Russia

Vertical Atoms, Chord Diagrams, and Knots (p.76) Stana Nikcevic, Mathematical Institute, SANU, Belgrade, Yugoslavia

Curvature Homogeneous Spaces and Related Topics (p.77) Marc Pauly, Université Libre de Bruxelles, Bruxelles, Belgium

Singular monopoles on compact 3-manifolds (p.77) Vyacheslav D. Sedykh, Gubkin Academy of Oil and Gas, Moscow, Russia

On classes of convex curves in projective spaces (p.80) Mamuka Shubladze, Georgian Technical University, Tbilisi, Georgia

The topology of hypersurface singularities with codimension one (p.80) Peter Topalov, Institute of Mathematics and Informatics, Sofia, Bulgaria

A family of 3-integrable metrics (p.83) Paweł Walczak, University of Lodz, Lodz, Poland

Foliations, finitely generated groups, and pseudogroups of transformations of manifolds carry some dynamics (p.83) Zuoling Zhou, Lingnan College of Zhongshan University, Guangzhou, People’s Republic of China

The Hausdorff measures of Sierpinski gasket and Koch curve (p.84)